OpenStudy (anonymous):

What is the midpoint between (-8, 5) and (2, -2)? (-5, 1.5) (-5, 3.5) (-6, 3) (-3, 1.5)

OpenStudy (anonymous):

\[\huge~\rm~m=\frac{ x_1+x_2 }{ 2 },\frac{ y_1+y_2 }{ 2 }\]

OpenStudy (anonymous):

\[\huge~\rm~m=\frac{ -8+2 }{ 2 },\frac{ 5-2}{ 2 }\]

OpenStudy (anonymous):

-3,1.5

OpenStudy (anonymous):

so answer D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you